OpenStudy (anonymous):

What operation must be done to both sides of the equation to solve p minus negative three and seven-ninths equals five-sixths? Add five-sixths Subtract three and seven-ninths Subtract five-sixths Add three and seven-ninths

OpenStudy (goldphenoix):

Hmm. The equation looks like: \[\large p - (-3) + \frac{ 7 }{ 9 } = \frac{ 5 }{ 6 }\]

OpenStudy (goldphenoix):

The multiple choice in my opinion doesn't really make sense, but the answer I would choose is, "Add three and seven-ninths"